# Quantum effects and temperature dependence of the free energy barrier and shear frequency in bilayer graphene

### v1

@@ Abstract
Little is known about the quantum effects of friction, and first-principles studies are virtually non-existent. In multilayer graphene, the order of the stacking of the layers plays a crucial role in the electronic properties and the manifestation of superconductivity, and friction affects how layers slide to change the stacking. Here, we study the effects of friction in bilayer graphene by focusing on quantum fluctuations in the free energy barrier that layers have to overcome to slide with respect to each other. We evaluate the free energy at the barrier in bilayer graphene through the self-consistent harmonic approximation, accounting for quantum and thermal ionic fluctuations. We obtain a very large reduction of the barrier of more than 30%. Temperature effects are weaker, although high temperatures induce a significant reduction of the barrier as well. Our approach is general and paves the way for systematically accounting for nuclear quantum and thermal effects in free energy barriers of other macroscopic systems.
@@
Superlubricity is another phenomenon present in systems involving graphene layers or graphite that has attracted a lot of attention(Liu et al., 2014; Sinclair et al., 2018), in which the energy landscape corresponding to the relative displacement of graphene layers is also central. Some works have studied how quantum effects, including zero-point renormalization (ZPR) and tunneling, affect the free-energy barrier of dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al. (2012); McIntosh et al. (2013); Cendagorta et al. (2016) of hydrogen. However, there are no studies of the quantum effects and temperature dependence of free-energy barriers in macroscopic systems.
@@
Another aspect of multilayer graphene and graphite that still needs further investigation and that serves as a consistency check on the adequacy of our approach is thermal expansion (TE). Ref. McQuade et al., 2021 measured it in bilayer and trilayer graphene for the first time. Ref. Mittal et al., 2021 uses QHA, which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972). Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al. (2018). However, in-plane PIMD results of Ref. Herrero and Ramírez, 2020 vastly overestimate the TE.
@@
Motivated by the aforementioned phenomena, especially by the lack of understanding of quantum effects in friction, we focus on studying the ZPR and temperature effects of the free-energy barrier. This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position. Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Monacelli et al. (2021). First, we compare our results to the height of the barrier with fixed atoms. Then, using also the quasiharmonic approximation (QHA), we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available. Finally, we determine the in-plane and out-of-plane TE coefficient in bilayer graphene and graphite, and compare our results to previous theoretical and experimental work.
@@
We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref. Libbi et al., 2020 and Ref. Wen et al., 2018, respectively. The out-of-plane interactions determine the barrier. The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig. 1, which gives good results when comparing to theory and experiment (see Table 1).
@@
Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical) between the interatomic potential we used in this workLibbi et al. (2020); Wen et al. (2018), LDA, and theoretical calculations and experiments from previous works. The values are a little bit lower but close to that of adiabatic-connection fluctuation-dissipation theorem within the random phase approximation (ACFDT-RPA)Zhou et al. (2015) and experimental dataLebedeva et al. (2017).
@@
The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$. For layers to slide with respect to each other, they have to overcome a barrier $V_{\mathrm{SP}}$. SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable. In order to determine the ZPR and temperature dependence of the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP.
The SSCHA minimizes the quantum free energy similarly to a variational Monte Carlo for nuclei, optimizing an ansatz density matrix of the most general Gaussian. The average atomic positions of the layers can be constrained at SP, and thanks to the simple shape of the variational ansatz, a quantum-free energy can be rigorously definedBianco et al. (2017); Monacelli et al. (2021). Notably, QHA cannot determine the barrier as the harmonic frequencies are not well-defined at SP. On the other hand, the SSCHA free energy landscape remains properly defined even in out-of-equilibrium and unstable configurations like SP. The Gaussian ansatz of the SSCHA allows separating the total free energy in an auxiliary harmonic contribution, which is long-ranged and can be interpolated up to a very fine $k$-mesh (40$\times$40), fundamental to avoid oversampling the energy of the shear mode, and a short-range anharmonic contribution, which can be properly converged already with small simulation cells. More details about QHA, SSCHA, and the interpolation method can be found in the Appendix.
@@
Figure 2: Barrier as a function of temperature. It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig. 1), to about 0.8 meV/atom, corresponding to a large ZPR of about 35%. The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K. The lines are a guide to the eye, as in the other figures of this work.
The barrier as a function of temperature can be seen in Fig. 2. The effects of the ZPR are very large: the barrier is reduced by an outstanding 35% with respect to the value at fixed nuclei. Since the frictional force is expected to be proportional to the barrier, quantum effects play a decisive role in the ability of layers to slide with respect to each other. Several other systems also present a reduction (sometimes also substantial) of the free energy relative to classical values, like adsorption or diffusion of hydrogen on several materialsMills and Jónsson (1994); Poma et al. (2012); McIntosh et al. (2013); Cendagorta et al. (2016). In terms of temperature effects, they are almost entirely covered by quantum fluctuations, since the barrier remains constant at around 0.8 meV/atom up to 500 K. Still when temperature increases up to 1000 K, the barrier is further reduced by 20% relative to the value at 100 K. Thus, high temperatures should help to reduce (static) friction between layers.
@@
At the equilibrium position, one of the vibrational modes at $\Gamma$ is the shear mode in both directions of the plane (doubly degenerate, see Fig. 3). Its temperature dependence can be determined using both SSCHA and QHA. In the QHA, the TE is obtained by determining the lattice parameters that minimize the free energy at each temperature. The harmonic phonon frequencies calculated at such lattice parameters then also give temperature dependent frequencies. See Appendix Sec. A1.1 for more details. The third out of phase mode of the layers is the breathing mode, in the out-of-plane direction. A discussion of its temperature dependence can also be found in the Appendix.
@@
The temperature dependence of the shear frequency can be seen in Fig. 4. The quantum effects evaluated within the SSCHA harden the shear frequency energy at low temperatures compared with the classical value at 0 K (dashed line). On the other hand, the ZPR in QHA softens the shear frequency. The temperature dependence of the shear mode is similar to the barrier, varying by about 20% from 0 to 1000 K, and having a larger dependence at temperatures above 500 K.
@@
In Fig. 5, we show the TE coefficient calculated with SSCHA and QHA in bulk, compared with experimental dataBailey and Yates (1970); Marsden et al. (2018) and PIMD calculations Herrero and Ramírez (2020). SSCHA results underestimate the thermal expansion. PIMD results are overall very good, although they also give values somewhat lower than experimentsBailey and Yates (1970); Marsden et al. (2018).
@@
To conclude, we demonstrated that quantum fluctuations are fundamental to determining properties of layered materials that account for friction. Using the self-consistent harmonic approximation, which can calculate the quantum free energy at unstable configurations, and a novel interpolation method, we unveiled the suppression of the barrier separating stable configurations in bilayer graphene. Quantum fluctuations reduce the free energy barrier by 35%, and gets further reduced at high temperatures. The shear modes, which occupy a small region of phase space, have to be adequately sampled to correctly determine the free energy barrier as well as the thermal expansion. Although our approach underestimates the out-of-plane thermal expansion compared with experimental values, in-plane agreement with experiments is excellent. We also obtained that the temperature dependence of the barrier is similar to that of the shear frequency, which agrees well with existing measurements.
@@ Acknowledgments
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 881603, and from the MORE-TEM ERC-SYN project, Grant Agreement No. 951215.
@@
A1 Calculation details
A1.1 Quasi-harmonic approximation (QHA)
$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right)\\ \end{split}$ (A1)
The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy. Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion. Using first-principles calculations, phonons can be determined using density functional perturbation theory (DFPT), which uses a primitive cell and is much faster than frozen phonon (finite differences) calculfations in supercells. However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature. For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter. Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see later Fig. A10), then an acoustic mode becomes unstable and the QHA is ill-defined. These issues are not present in SSCHA.
A1.2 Stochastic self-consistent harmonic approximation (SSCHA)
In general, the free-energy of a ionic Hamiltonian $H=T+V$ is given by
$F_{H}=\mathrm{tr}(\rho_{H}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{H}\mathrm{ln}\rho_{H})$ (A2)
$\mathcal{F}_{H}(\mathcal{H})=\mathrm{tr}(\rho_{\mathcal{H}}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{\mathcal{H}}\mathrm{ln}\rho_{\mathcal{H}})$ (A3)
[…]
where $\mathbf{R}$ indicates a general ionic configuration and $\rho_{\mathcal{H}}(\mathbf{R})=\langle\mathbf{R}|e^{-\beta\mathcal{H}}|\mathbf{R}\rangle/Z_{\mathcal{H}}$ is the probability density of finding the crystal in a generic configuration $\mathbf{R}$ ($Z_{\mathcal{H}}=\mathrm{tr}[e^{-\beta\mathcal{H}}]$ the partition function). By stochastically sampling configurations $\mathbf{R}$ for a given density matrix $\rho_{\mathcal{H}}$, and determining the gradient of the free energy with respect the harmonic parameters (centroids and force constant matrix elements), the free energy can be minimized self-consistently. We used 24000 configurations in all calculations. A detailed description of SSCHA can be found for example in Refs. Errea et al., 2014 or Monacelli et al., 2021.
The logarithm term in Eq. A1 can be hard to converge for low frequencies. So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way, and $F_{\mathcal{H}}$ can be calculated using a dense grid. This is the novel SSCHA interpolation method we mentioned in the main section, which was used with a $40\times 40$ in-plane grid to determine the barrier and also the out-of-plane TE. To properly determine the barrier, the interpolation scheme is used both at equilibrium and at SP, and then the interpolated free energies are subtracted.
The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy. In a purely harmonic system, such frequencies correspond to the physical frequencies. But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al. (2021). Later in Fig. A5, we show the temperature dependendence of both the auxiliary and physical frequencies.
A2 Barrier
@@
To determine the barrier, the free energy has to be determined at the equilibrium position and SP. At the equilibrium position, there are two shear modes, indicated with $\mathbf{S}$ in Fig. 3(b). As mentioned in the main text, the mode perpendicular to the bond direction shown in Fig. 1 is stable in SP, while the shear mode along the reaction coordinate (using the language of chemical reactions) is unstable. As opposed to the breathing mode, that is flat around $\Gamma$, the shear modes are only present in a small region of the BZ ($1/N^{2}$, with $N$ larger than 20), as a result of the large speed of sound in graphene. So in small supercells, the shear modes will be over-represented. This is more problematic at SP, where the SSCHA frequency corresponding to she unstable mode takes very small values (see Fig. A1), which makes the convergence of the free energy hard, mostly due to the logarithm term in $F_{\mathcal{H}}$ in Eq. (A4). Just determining the barrier with this equation and a $N\times N\times 1$ supercell, results in Fig. A2, which is not well converged.
@@
In order to get better convergence, we use the interpolation method we mentioned in Sec. A1.2, and obtain a much more stable barrier as a function of $1/N$ (see Fig. A3). The $40\times 40$ interpolation grid used for $F_{\mathcal{H}}$ corresponds to a supercell of 6400 atoms. For 100, 200 and 300 K, the barrier seems well converged. For a 1000 K, the trend seems to indicate convergence within 0.02 meV/atom, so for the intermediate temperatures 400 and 500 K, the barrier is likely also converged within a similar error, resulting in Fig. 2 of the main text.
@@
Figure A2: Barrier as a function of temperature, for different supercells, just using the $N\times N\times 1$ grid to determine the harmonic part of the free energy Eq. (A4). The barrier changes significantly with $N$, which means the shear mode is overrepresented. On the other hand, interpolating the harmonic part of the free energy, convergence is much better. See Fig. A3.
Figure A3: Barrier as a function of $1/N$, for different temperatures, interpolating the first term of Eq. A4. The barrier seems well converged at low temperatures and harder to converge at higher temperatures. Values are likely converged within less than 0.05 meV/atom.
@@
A3 Shear and breathing frequency
@@
A4 Thermal expansion (TE)
The interlayer distance and in-plane lattice parameter obtained in bulk with SSCHA can be seen in Fig. A8. The fit corresponds to Eq. (6) of Ref. Zorzi and Perottoni, 2021, both for in and out-of-plane TE ($b$ can be $a$ or $c_{d}$),
$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (A5)
which depends on parameters $b_{0}$, $X$ and $\theta$. In Fig. A9 we compare SSCHA and QHA for bilayer graphene and bulk. The TE for both SSCHA and QHA bulk is lower relative to bilayer. In-plane, Fig. A10, we can also see a clear reduction of the TE coefficient from monolayer to bulk. In what follows, we describe in detail convergence studies we did for SSCHA, and also for QHA calculations.
Figure A8: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (A5).
@@
In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (A15).
@@
On the other hand, $a(T)$ is already converged at about $N=8$ (Fig. A17).
A4.2 Bulk
@@
SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. A8 a.
In-plane, based on $N=8$ already giving converged results in bilayer (Fig. A17), and also considering that the QHA convergence in bulk (Fig. A23) is much faster than in bilayer (Fig. A15), $N=8$ should give converged SSCHA results in bulk, which corresponds to the blue dots in Fig. A8 b.

### v2

@@ Abstract
Little is known about the quantum effects of friction, and first-principles studies are virtually non-existent. In multilayer graphene, the order of the stacking of the layers plays a crucial role in the electronic properties and the manifestation of superconductivity, and friction affects how layers slide to change the stacking. Here, we study the effects of friction in bilayer graphene by focusing on quantum fluctuations in the free energy barrier that layers have to overcome to slide with respect to each other. We evaluate the free energy at the barrier in bilayer graphene through the self-consistent harmonic approximation, accounting for quantum and thermal ionic fluctuations, and obtain a very large reduction of the barrier of more than 30% at low temperatures. There is little temperature dependence up to about 500 K, but it increases at higher temperatures. Our approach is general and paves the way for systematically accounting for nuclear quantum and thermal effects in free energy barriers of other macroscopic systems.
I Introduction
@@
Superlubricity is another phenomenon present in systems involving graphene layers or graphite that has attracted a lot of attention(Liu et al., 2014; Sinclair et al., 2018), in which the energy landscape corresponding to the relative displacement of graphene layers is also central. Yet another phenomena involving sliding surfaces is thermolubricity, which refers to the assistance from thermal excitations to overcome barriers. This term has been coined by some works that have studied the transition from stick-slip motion into stochastic fluctuations in friction at sufficiently high temperaturesBaykara et al. (2018). A few works have studied how quantum effects, including zero-point renormalization (ZPR) and tunneling, affect the free-energy barrier of dissociative adsorptionMills and Jónsson (1994) and diffusionPoma et al. (2012); McIntosh et al. (2013); Cendagorta et al. (2016) of hydrogen. However, there are no studies of the quantum effects and temperature dependence of free-energy barriers in macroscopic systems.
@@
Another aspect of multilayer graphene and graphite that still needs further investigation and that serves as a consistency check on the adequacy of our approach is thermal expansion (TE). Ref. McQuade et al., 2021 measured it in bilayer and trilayer graphene for the first time. Ref. Mittal et al., 2021 uses the quasi-harmonic approxmimation (QHA), which underestimates the out-of-plane TE at low temperaturesBailey and Yates (1970) and has a very rapid increase at high temperatures; in-plane, it also underestimates the TE at low temperaturesBailey and Yates (1970) and overestimates it at high temperaturesMorgan (1972). Much better out-of-plane results were obtained by using path integral molecular dynamics (PIMD)Herrero and Ramírez (2020), but while the TE flattens at 500 K in graphite, experimental values continue increasingMorgan (1972); Marsden et al. (2018). However, in-plane PIMD results of Ref. Herrero and Ramírez, 2020 vastly overestimate the TE.
@@
Motivated by the aforementioned phenomena, especially by the lack of understanding of quantum effects in friction, we focus on studying the ZPR and temperature effects of the free-energy barrier. This is a very challenging problem due to the difficulty of determining the free energy in a non-equilibrium position. Here we tackle it using the state-of-the-art stochastic self-consistent harmonic approximation (SSCHA)Monacelli et al. (2021). First, in Sec.II, we briefly introduce SSCHA and the QHA, and explain how is the barrier determined. Then, in Sec. III, we present our main results and compare to the height of the barrier with fixed atoms. Subsequently, using also the QHA, we calculate the temperature dependence of the shear frequency and compare to the limited experimental data available. Finally, we determine the in-plane and out-of-plane TE coefficient in bilayer graphene and graphite, and compare our results to previous theoretical and experimental work. In Sec. IV, we summarize the conclusions of our work.
II Theoretical and Computational Framework
Here, we briefly describe the QHA, SSCHA, and a new SSCHA interpolation method used to determine the barrier. We also introduce the equation that is used to fit SSCHA thermal expansion results. A more detailed description of SSCHA can be found in Refs. Errea et al., 2014 or Monacelli et al., 2021.
II.1 Quasi-harmonic approximation (QHA)
$\begin{split}F(T,\{a_{i}\})=&U_{\mathrm{lat}}(\{a_{i}\})+\frac{1}{N}\sum_{\mathbf{q}\nu}\frac{1}{2}\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})+\\ &+\frac{1}{N}\sum_{\mathbf{q}\nu}kT\mathrm{ln}\left(1-e^{-\hbar\omega_{\mathbf{q}\nu}(\{a_{i}\})/kT}\right).\\ \end{split}$ (1)
The value of the lattice parameters at a given temperature, $a_{i}(T)$, is determined by minimizing the free energy. Using an interatomic potential, the phonon frequencies at each value of the lattice parameters can be determined with little computational effort, and thus the QHA offers a quick method to determine the lattice expansion. Using first-principles calculations, phonons can be determined using density functional perturbation theory (DFPT), which uses a primitive cell and is much faster than frozen phonon (finite differences) calculations in supercells. However, it does not account for the “true anharmonicity”Allen (2019), present at fixed values of the parameters when varying temperature. For example, it gives an increasing as opposed to decreasing temperature dependence of the G mode in graphene, due to the negative TE of the lattice parameter. Also, if the TE is negative and the lattice parameter reduces below the classical lattice value, as in the case of monolayer graphene at about 400 K (see later Fig. A10), then an acoustic mode becomes unstable and the QHA is ill-defined. These issues are not present in SSCHA.
II.2 Stochastic self-consistent harmonic approximation (SSCHA)
In general, the free-energy of a ionic Hamiltonian $H=T+V$ is given by
$F_{H}=\mathrm{tr}(\rho_{H}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{H}\mathrm{ln}\rho_{H}),$ (2)
$\mathcal{F}_{H}(\mathcal{H})=\mathrm{tr}(\rho_{\mathcal{H}}H)+\frac{1}{\beta}\mathrm{tr}(\rho_{\mathcal{H}}\mathrm{ln}\rho_{\mathcal{H}}).$ (3)
[…]
where $\mathbf{R}$ indicates a general ionic configuration and $\rho_{\mathcal{H}}(\mathbf{R})=\langle\mathbf{R}|e^{-\beta\mathcal{H}}|\mathbf{R}\rangle/Z_{\mathcal{H}}$ is the probability density of finding the crystal in a generic configuration $\mathbf{R}$ ($Z_{\mathcal{H}}=\mathrm{tr}[e^{-\beta\mathcal{H}}]$ is the partition function). By stochastically sampling configurations $\mathbf{R}$ for a given density matrix $\rho_{\mathcal{H}}$, and determining the gradient of the free energy with respect the harmonic parameters (centroids and force constant matrix elements), the free energy can be minimized self-consistently. Thanks to the variational ansatz, the SSCHA quantum free energy is rigorously defined, even in out-of-equilibrium and unstable configurations like SP. Notably, on the other hand, the QHA cannot determine the barrier, since the harmonic frequencies are not well defined at SP.
The logarithm term in the harmonic term Eq. (1) can be hard to converge for low frequencies. Separating the total free energy as in Eq. (4) allows to interpolate the auxiliary harmonic term up to a very fine $q$-mesh, which is fundamental to avoid oversampling the energy of the long-ranged shear modes, while the anharmonic contribution can be properly converged already with small simulation cells. So rather than calculating $F_{\mathcal{H}}$ directly with the grid corresponding to the SC used in the minimization procedure, the force constant matrix can be Fourier interpolated in the usual way, and $F_{\mathcal{H}}$ can be calculated using a dense mesh.
This procedure is followed both at SP and the equilibrium position, and the barrier is obtained by taking the difference of the corresponding free energies. More technical details are provided in the Appendix, Sec. A1.
The SSCHA frequencies are in principle auxiliary frequencies that minimize the free energy. In a purely harmonic system, such frequencies correspond to the physical frequencies. But in general, they have to be corrected with the so-called “bubble correction” (and also an additional higher order term in very anharmonic cases)Monacelli et al. (2021), which we do in Sec. III.2 to obtain the temperature dependence of the shear frequency. Convergence with the supercell size of both the physical and auxiliary frequencies can be seen in the Appendix, Fig. A5.
II.3 Thermal expansion (TE)
As opposed to the barrier, the thermal expansion (TE) can also be calculated using the QHA, which although is much faster, is also in principle less accurate. With both QHA and SSCHA we determine the in-plane and out-of-plane lattice parameters as a function of temperature, $a(T)$ and $c(T)$, to then determine the in-plane and out-of-plane coefficients of TE, $\alpha_{\mathrm{a}}(T)=\tfrac{1}{a}\tfrac{da}{dT}$ and $\alpha_{\mathrm{c}}(T)=\tfrac{1}{c}\tfrac{dc}{dT}$. In the QHA case, the TE can calculated quickly and accurately for many temperatures, so we just use finite differences. For SSCHA, to reduce numerical error, we use Eq. (6) of Ref. Zorzi and Perottoni, 2021,
$b(T)=b_{0}\mathrm{exp}\left(\frac{X\theta}{e^{X/T}-1}\right),$ (5)
which depends on parameters $b_{0}$, $X$ and $\theta$, to fit $a(T)$ and $c_{d}(T)$ ($b$ can be $a$ or $c_{d}$). Then we determine the coefficients of TE from the derivative of the fit.
III Results
III.1 Barrier
@@
We first focus on studying bilayer graphene, where the in-plane and (most importantly) out-of-plane interactions are given by the interatomic potential of Ref. Libbi et al., 2020 and Ref. Wen et al., 2018, respectively. The out-of-plane interactions determine the barrier. The energy profile of a layer moving with respect to the other one along the bond direction is given by Fig. 1, which gives good reults when comparing to theory and experiment (see Table 1).
@@
Table 1: Comparison of the barrier $V_{\mathrm{SP}}$ and the shear frequency at fixed ionic positions (classical at $T=0$ K) between the interatomic potential we used in this workLibbi et al. (2020); Wen et al. (2018), LDA, and theoretical calculations and experiments from previous works. The values are a little bit lower but close to that of adiabatic-connection fluctuation-dissipation theorem within the random phase approximation (ACFDT-RPA)Zhou et al. (2015) and experimental dataLebedeva et al. (2017).
@@
The lower layer is fixed in position A, and there are two minima of the same energy, AB and AC, at a distance of a bond length $d=1.42$. For layers to slide with respect to each other, they have to overcome a barrier $V_{\mathrm{SP}}$. SP is a saddle point in the full energy landscape, so the shear mode in the perpendicular direction (along the direction of two nearby AA maxima) is stable. In order to determine the ZPR and temperature dependence of the barrier, we use SSCHA to determine the free energy at the equilibrium position and SP, using the interpolation method described in Sec. II.2.
@@
Figure 2: Barrier as a function of temperature. It is reduced from 1.24 meV/atom, the value at fixed nuclei (VSP in Fig. 1), to about 0.8 meV/atom, corresponding to a large reduction of about 35% due to anharmonicities. The barrier does not change much up to about 500 K, but the value at 1000 K is about 20% smaller relative to the value at 100 K. The lines are a guide to the eye, as in the other figures of this work.
[…]
To confirm that quantum fluctuations are very minor down to 100 K, we performed SSCHA calculations in the classical limit $\hbar\omega/kT\rightarrow 0$ (which recovers the classical Boltzmann distribution for the distorted configurations) at 100 K and 300 K. We obtained comparable results, except for a small difference of less than 0.05 meV/atom at 100 K with respect to the quantum values. Much lower temperatures would have to be considered to recover the classical value at fixed ions. Thus, it is the anharmonic vibrations that play a decisive role in the ability of layers to slide with respect to each other.
In a previous workNery et al. (2020), using density functional theory with an LDA functional (see Table 1), some of us obtained that the maximum shear stress in simulated stick-slip motion is 0.2 GPa, and we mentioned a good agreement with experimental values 0.14 GPa (Ref. Liu et al., 2012) and 0.1 GPa (Ref. Dienwiebel et al., 2004). Assuming the maximum shear stress decreases by 35% just like the barrier, we obtain 0.13 GPa, in excellent agreement with experiments.
Other systems involving hydrogen diffusion across a free energy barrier do present significant differences between classical and quantum free energy barriers due to the large quantum fluctuations of hydrogen. Such differences have been observed at 100 K (Ref. Mills and Jónsson, 1994), 250 K (Ref. McIntosh et al., 2013), and even higher temperaturesPoma et al. (2012), although other worksCendagorta et al. (2016) obtain very similar barriers at already 200 K.
Regarding the evolution of the barrier as the temperature increases, there is a weak dependence up to about 500 K, since the barrier remains close to around 0.8 meV/atom. However, when temperature increases up to 1000 K, the barrier is further reduced by 20% relative to the value at 100 K. Thus, high temperatures should help to reduce (static) friction between layers. In particular, it could aid the production of ABC-stacked multilayer graphene from its more common AB-stacked version through shear stress, as mentioned in the Introduction. More in general, high temperatures could also help engineer specific stacking arrangements in other layered materials.
III.2 Shear frequency
@@
At the equilibrium position, one of the vibrational modes at $\Gamma$ is the shear mode in both directions of the plane (doubly degenerate, see Fig. 3). Its temperature dependence can be determined using both SSCHA and QHA. In the QHA, the TE is obtained by determining the lattice parameters that minimize the free energy at each temperature.
The third out of phase mode of the layers is the breathing mode, in the out-of-plane direction. A discussion of its temperature dependence can be found in the Appendix.
@@
The temperature dependence of the shear frequency can be seen in Fig. 4. The quantum effects evaluated within the SSCHA harden the shear frequency energy at low temperatures compared with the classical value at 0 K (dashed line). On the other hand, the ZPR in QHA softens the shear frequency. The temperature dependence of the shear mode is similar to that of the barrier, varying by about 20% from 0 to 1000 K, and having a larger dependence at temperatures above 500 K.
@@ III.3 Thermal expansion
In Fig. 5, we show the TE coefficient calculated with SSCHA and QHA in bulk, compared with experimental dataBailey and Yates (1970); Marsden et al. (2018) and PIMD calculations Herrero and Ramírez (2020). SSCHA curves result from fitting Eq. (5) to $a(T)$ and $c(T)$ (see Fig. A8) and then taking the derivative. SSCHA results underestimate the thermal expansion. PIMD results are overall very good, although they also give values somewhat lower than experimentsBailey and Yates (1970); Marsden et al. (2018).
@@
IV Conclusions
To conclude, we demonstrated that anharmonicities are fundamental to determining properties of layered materials that account for friction. Using the self-consistent harmonic approximation, which can calculate the quantum free energy at unstable configurations, and a novel interpolation method, we unveiled the suppression of the barrier separating stable configurations in bilayer graphene. Anharmonicities reduce the free energy barrier by 35%, and gets further reduced at high temperatures. On the other hand, quantum fluctuations are not important down to 100 K. The shear modes, which occupy a small region of phase space, have to be adequately sampled to correctly determine the free energy barrier as well as the thermal expansion. Although our approach underestimates the out-of-plane thermal expansion compared with experimental values, in-plane agreement with experiments is excellent. We also obtained that the temperature dependence of the barrier is similar to that of the shear frequency, which agrees well with existing measurements.
@@ Acknowledgments
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No. 881603, and from the MORE-TEM ERC-SYN project, Grant Agreement No. 951215. J.P.N. is currently supported by the European Union under a Marie Skłodowska-Curie Individual Fellowship, Project GreenNP No. 101151380.
@@
A1 Barrier
@@
To determine the barrier, the free energy has to be determined at the equilibrium position and SP. At the equilibrium position, there are two shear modes, indicated with $\mathbf{S}$ in Fig. 3(b). The mode perpendicular to the bond direction shown in Fig. 1 is stable in SP, while the shear mode along the reaction coordinate (using the language of chemical reactions) is unstable. As opposed to the breathing mode, that is flat around $\Gamma$, the shear modes are only present in a small region of the BZ ($1/N^{2}$, with $N$ larger than 20), as a result of the large speed of sound in graphene. So in small supercells, the shear modes will be over-represented. This is more problematic at SP, where the SSCHA frequency corresponding to she unstable mode takes very small values (see Fig. A1), which makes the convergence of the free energy hard, mostly due to the logarithm term in $F_{\mathcal{H}}$ in Eq. (4). Just determining the barrier with this equation and a $N\times N\times 1$ supercell, results in Fig. A2, which is not well converged.
@@
In order to get better convergence, we use the interpolation method mentioned in Sec. II.2, and obtain a much more stable barrier as a function of $1/N$ (see Fig. A3). The $40\times 40$ interpolation grid used for $F_{\mathcal{H}}$ corresponds to a supercell of 6400 atoms. For 100, 200 and 300 K, the barrier seems well converged. For a 1000 K, the trend seems to indicate convergence within 0.02 meV/atom, so for the intermediate temperatures 400 and 500 K, the barrier is likely also converged within a similar error, resulting in Fig. 2 of the main text.
@@
Figure A2: Barrier as a function of temperature, for different supercells, just using the $N\times N\times 1$ grid to determine the harmonic part of the free energy Eq. (4). The barrier changes significantly with $N$, which means the shear mode is overrepresented. On the other hand, interpolating the harmonic part of the free energy, convergence is much better. See Fig. A3.
Figure A3: Barrier as a function of $1/N$, for different temperatures, interpolating the first term of Eq. 4. The barrier seems well converged at low temperatures and harder to converge at higher temperatures. Values are likely converged within less than 0.05 meV/atom.
@@
A2 Shear and breathing frequency
@@
A3 Thermal expansion (TE)
For the SSCHA out-of-plane TE calculations, we also used the interpolation scheme with a $40\times 40$ grid. The interlayer distance and in-plane lattice parameter obtained in bulk with SSCHA can be seen in Fig. A8. The derivative of these fits results in the SSCHA curves of Fig. 5.
In Fig. A9 we compare SSCHA and QHA for bilayer graphene and bulk. The TE for both SSCHA and QHA bulk is lower relative to bilayer. In-plane, Fig. A10, we can also see a clear reduction of the TE coefficient from monolayer to bulk. In what follows, we describe in detail convergence studies we did for SSCHA, and also for QHA calculations.
Figure A8: Out-of plane (a) and in-plane (b) SSCHA thermal expansion (blue dots). The fit corresponds to Eq. (5).
@@
In-plane, a larger $N$ of about 50 is needed to fully achieve convergence (Fig. A15).
@@
On the other hand, $a(T)$ has less variation with temperature relative to QHA and is already converged at about $N=8$ (Fig. A17).
A3.2 Bulk
@@
SSCHA. Here we also used the interpolation method ($40\times 40$ grid), to go beyond the $N=8$ supercell that was possible in the direct SSCHA calculation, resulting in the blue dots of Fig. A8 (a).
In-plane, based on $N=8$ already giving converged results in bilayer (Fig. A17), and also considering that the QHA convergence in bulk (Fig. A23) is much faster than in bilayer (Fig. A15), $N=8$ should give converged SSCHA results in bulk, which corresponds to the blue dots in Fig. A8 (b).